1) At a particular temperature k= 6.7x10^-6 for the following reaction:

2Br2(g) +O2(g) -> 2Br2O(g)

In an experiment in which 1.0 mol of Br2 and 1.0 mol of O2 are mixed in a 2.5-L vessel, what are the equilibrium concentrations of all gases?

2) At a particular temperature k= 6.5x10^-6 for the following reaction:

2Cl2(g)+O2(g) -> 2Cl2O(g)

In an experiment in which 2.0 mol of Cl2 and 1.5 mol of O2 are mixed in a 2.0-L vessel, what are the equilibrium concentrations of all gases?


1
Expert's answer
2022-02-14T17:47:04-0500

1) 2Br2(g) + O2(g) -> 2Br2O(g)

k = [Br2O]2/[O2][Br2]2

C0(Br2O) = 0 mol/L

C0(O2) = 1 mol / 2.5 L = 0.4 mol/L

C0(Br2) = 1 mol / 2.5 L = 0.4 mol/L

[Br2O] = 2x mol/L

[O2] = (0.4 - x) mol/L

[Br2] = (0.4 - 2x) mol/L

6.7 x 10-6 = (2x)2/(0.4 - x)(0.4 - 2x)2

x = 0.52 x 10-3 mol/L

[Br2O] = 0.001 mol/ L

[O2] = 0.399 mol/L

[Br2] = 0.399 mol/L

2) 2Cl2(g) + O2(g) -> 2Cl2O(g)

k = [Cl2O]2/[O2][Cl2]2

C0(Cl2O) = 0 mol/L

C0(O2) = 1.5 mol / 2.0 L = 0.74 mol/L

C0(Cl2) = 2.0 mol / 2.0 L = 1.0 mol/L

[Cl2O] = 2x mol

[O2] = (0.74 - x) mol/L

[Cl2] = (1.0 - 2x) mol/L

6.5 x 10-6 = (2x)2/(0.74 - x)(1.0 - 2x)2

x = 1.1 x 10-3 mol/L

[Cl2O] = 0.0022 mol/ L

[O2] = 0.739 mol/L

[Cl2] = 0.999 mol/L